# A proof of the Kim-Vu sandwich conjecture

### v1

@@ 9.1 Translating likely properties of $G_{d}(n)$ into subgraphs of $F(n,d,m)$
$\eta=\frac{C_{0}}{40}\max\left\{\frac{\mu}{\log n},\left(\frac{n\log n}{m}\right)^{1/8}\right\}.$ (53)
@@
Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$. Let $E$ be a set of $m$ edges chosen uniformly at random from $n]^{(2)}\setminus E(K)$, and let $F=K+E$. Furthermore, form $E^{\prime}$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F^{\prime}=K+E^{\prime}$. Note that $p$ is chosen so that $\mathbb{P}(|E^{\prime}|=i)$ is maximised at $i=m$, and, thus $\mathbb{P}(|E^{\prime}|=m)\geq 1/n^{2}$. We will show that, that with probability $1-o(n^{-10})$, LABEL:prop:degreebound, LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating hold for $F^{\prime}$ in place of $F$. From this, it holds that, with probability $1-o(n^{-8})$, LABEL:prop:degreebound, LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating hold for $F$. Then, by Lemma 9.1, with probability $1-o(n^{-2})$ we have that LABEL:prop:degreebound holds and, for all but at most $n^{-6}|\mathcal{K}_{d}(F)|$ graphs $K\in\mathcal{K}_{d}(F)$ we have that LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating hold. Therefore, by Theorem 4.1, with probability $1-o(n^{-2})$, (6) holds for each $e,f\notin E(F)$.
@@ 10 The lower part of the sandwich
In this section, we give the lower part $(G_{*},G)$ of the sandwich and prove it satisfies Theorem 1.1. As noted, this part of the theorem previously follows from the combination of work by Kim and Vu [20], Dudek, Frieze, Ruciński and Šileikis [9] and Gao, Isaev and McKay [14, 15]. Moreover, in [13], Gao, Isaev and McKay use an edge addition process that is effectively the same as we use here, and analyse it as long as $d=\omega(\log n)$. We include the lower part of the sandwich for completion, and because, due to the use of our corresponding version of Lemma 9.1, the analysis is different. For the coupling $(G_{*},G)$, the process used to create $G$ is the complement of the process used for the upper part, while for ease of analysis we generate $G_{*}$ via a sequence of graphs $H_{0},H_{1},\ldots$ chosen to be most like the complement of the graphs $G_{0}^{+},G_{1}^{+},\ldots$ in Section 3.5. This coupling is given in Section 10.1. We then given an overview of the analysis of the process required in Section 10.2, and outline its proof over the rest of this section.
@@ 10.2 Analysis overview
Now, similarly to $H_{i}$ in Section 3.5, $e(H_{N})\sim\mathrm{Bin}(N,1-\eta)$ so, by a Chernoff bound we have that $e(H_{N})\geq(1-2\eta)N$ with high probability. As, by another Chernoff bound, $M\leq(1-\varepsilon+\varepsilon/2)\binom{n}{2}$ with high probability, we have $e(H_{N})\geq M$ with high probability. Thus, to prove $G_{*}\subset G$ with high probability it is sufficient to prove that $H_{N}\subset G$ with high probability.
Similarly to Claim 3.1 b), we have the following claim, which we prove in Section 10.3. Recall that we say $F\sim F^{-}(n,d,m)$ if $F$ is a random graph with vertex set $[n]$ which has the same distribution as HEH-E if $H\sim G_{d}(n)$ and $E$ is a uniformly-chosen set of $m$ edges from $E(H)$.
@@
$|\{K\in\mathcal{K}_{d}(n):F_{i-1}+e\subset E(K)\}|\leq\left(1+\eta\right)|\{K\in\mathcal{K}_{d}(n):F_{i-1}+f\subset E(K)\}|.$ (54)
Suppose, then, that $e\in E(H_{N})$. Thus, there is some $1\leq i\leq N$ with $e_{k(i)}=e$, with $e$ not in $e_{1},\dots,e_{k(i)-1}$, and with $x_{k(i)}\leq 1-\eta$. Now, the edges $e_{\ell(1)},e_{\ell(2)},\ldots,e_{\ell(N)}$ are all distinct, so we must have $\ell(N)\geq k(N)$. Thus, there is some $j\in[N]$ with $\ell(j-1)<k(i)\leq\ell(j)$. As $e$ is not in $e_{1},\dots,e_{k(i)-1}$, and (54) holds, we therefore have $\ell(j)=k(i)$ and $e=e_{\ell(j)}\in E(F_{j})\subset E(F_{dn/2})=E(F)$. Thus, $H_{N}\subset F$.
@@ 10.4 Concentration for switching paths
Let $1/n\ll 1/C\ll\eta,\varepsilon$. Let $d\geq C\log n$ and $\varepsilon dn\leq m\leq dn/2$. Let $K\sim K_{d}(n)$, let $E$ be a uniformly random set of $m$ edges from $E(K)$, and let F=KEF=K-E. Then, with probability $1-o(n^{-8})$, the following holds.
@@
$|\hat{E}|\geq(n-1-d)n-2(n-1-d)(13+2(n-1-d))\geq\left(1-\frac{\eta}{10}\right)(n-1-d)n,$
so that $|\hat{E}|=\left(1\pm\frac{\eta}{10}\right)(n-1-d)n$.
@@
Therefore, with probability $1-o(n^{-21})$, the number of $(K\setminus F,K_{n}\setminus K)$-alternating $v_{1},y$-paths with length 3 avoiding $Z\cup\{x\}$ is, as $d=\left(1\pm\frac{\eta}{10}\right)n$,

### v2

@@ 9.1 Translating likely properties of $G_{d}(n)$ into subgraphs of $F(n,d,m)$
$\eta=C_{0}\max\left\{\frac{\mu}{\log n},\left(\frac{n\log n}{m}\right)^{1/8}\right\}.$ (54)
@@
Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$. Let $E$ be a set of $m$ edges chosen uniformly at random from $n]^{(2)}\setminus E(K)$, and let $F=K+E$. Furthermore, form $E^{\prime}$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F^{\prime}=K+E^{\prime}$. Note that $p$ is chosen so that $\mathbb{P}(|E^{\prime}|=i)$ is maximised at $i=m$, and, thus $\mathbb{P}(|E^{\prime}|=m)\geq 1/n^{2}$. We will show that, that with probability $1-o(n^{-10})$, LABEL:prop:degreebound, LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating hold for $F^{\prime}$ in place of $F$. From this, it holds that, with probability $1-o(n^{-8})$, LABEL:prop:degreebound, LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating (defined in Theorem 4.1) hold for $F$. Then, by Lemma 9.1, with probability $1-o(n^{-2})$ we have that LABEL:prop:degreebound holds and, for all but at most $n^{-6}|\mathcal{K}_{d}(F)|$ graphs $K\in\mathcal{K}_{d}(F)$ we have that LABEL:prop:lowerboundKFKalternating and LABEL:prop:upperboundFKKalternating hold. Therefore, by Theorem 4.1, with probability $1-o(n^{-2})$, (6) holds for each $e,f\notin E(F)$.
@@ 10 The lower part of the sandwich
In this section, we give the lower part $(G_{*},G)$ of the sandwich and prove it satisfies Theorem 1.1. As noted, this part of the theorem previously follows from the combination of work by Kim and Vu [20], Dudek, Frieze, Ruciński and Šileikis [9] and Gao, Isaev and McKay [14, 15]. Moreover, in [13], Gao, Isaev and McKay use an edge addition process that is effectively the same as we use here, and analyse it as long as $d=\omega(\log n)$. We include the lower part of the sandwich for completion, and because, due to the use of our corresponding version of Lemma 9.1, the analysis is different. For the coupling $(G_{*},G)$, the process used to create $G$ is the complement of the process used for the upper part, while for ease of analysis we generate $G_{*}$ via a sequence of graphs $H_{0},H_{1},\ldots$ chosen to be more like the complement of the graphs $G_{0}^{+},G_{1}^{+},\ldots$ in Section 3.5. This coupling is given in Section 10.1. We then given an overview of the analysis of the process required in Section 10.2, and outline its proof over the rest of this section.
@@ 10.2 Analysis overview
Now, similarly to $H_{i}$ in Section 3.5, $e(H_{N})\sim\mathrm{Bin}(N,1-\eta)$ so, by a Chernoff bound we have that $e(H_{N})\geq(1-2\eta)N\geq(1-3\eta)dn/2$ with high probability. As, by another Chernoff bound, $M\leq(1-\varepsilon+\varepsilon/2)dn/2$ with high probability, we have $e(H_{N})\geq M$ with high probability. Thus, to prove $G_{*}\subset G$ with high probability it is sufficient to prove that $H_{N}\subset G$ with high probability.
Similarly to Claim 3.1 b), we have the following claim, which we prove in Section 10.3. Recall from Definition 2.1 that we say $F\sim F^{-}(n,d,m)$ if $F$ is a random graph with vertex set $[n]$ which has the same distribution as HEH-E where $H\sim G_{d}(n)$ and $E$ is a uniformly-chosen set of $m$ edges from $E(H)$.
@@
$|\{K\in\mathcal{K}_{d}(n):F_{i-1}+e\subset E(K)\}|\leq\left(1+\eta\right)|\{K\in\mathcal{K}_{d}(n):F_{i-1}+f\subset E(K)\}|$
and in particular,
$\frac{|\{K\in\mathcal{K}_{d}(n):F_{i-1}+e_{\ell(i)}\subset K\}|}{\max_{e\notin E(F_{i-1})}|\{K\in\mathcal{K}_{d}(n):F_{i-1}+e\subset K\}|}\geq\frac{1}{1+\eta}>1-\eta.$ (55)
Suppose, then, that $e\in E(H_{N})$. Thus, there is some $1\leq i\leq N$ with $e_{k(i)}=e$, with $e$ not in $e_{1},\dots,e_{k(i)-1}$, and with $x_{k(i)}\leq 1-\eta$. Now, the edges $e_{\ell(1)},e_{\ell(2)},\ldots,e_{\ell(N)}$ are all distinct, so we must have $\ell(N)\geq k(N)$. Thus, there is some $j\in[N]$ with $\ell(j-1)<k(i)\leq\ell(j)$. As $e$ is not in $e_{1},\dots,e_{k(i)-1}$, $x_{k(i)}\leq 1-\eta$ and (55) holds, we therefore have $\ell(j)=k(i)$ and $e=e_{\ell(j)}\in E(F_{j})\subset E(F_{dn/2})=E(F)$. Thus, $H_{N}\subset F$.
@@ 10.4 Concentration for switching paths
Let $1/n\ll 1/C\ll\eta,\varepsilon$. Let $d\geq C\log n$ and $\varepsilon dn\leq m\leq dn/2$. Let $K\sim G_{d}(n)$, let $E$ be a uniformly random set of $m$ edges from $E(K)$, and let F=KEF=K-E. Then, with probability $1-o(n^{-8})$, the following holds.
@@
$|\hat{E}|\geq(n-1-d)n-2(n-1-d)(13+2(n-1-d))\geq\left(1-\frac{\eta}{10}\right)(n-1-d)n,$ (56)
@@
Let $p=2m/dn$. Let $E^{\prime}\subset E(K)$ be chosen by including each edge in $K$ independently at random with probability $p$, so that $\mathbb{P}(|E^{\prime}|=m)\geq 1/n^{2}$, and let $F^{\prime}=K-E^{\prime}$. For each $i\in[n-d+1]$, let $X_{i}=|\hat{E}_{i}\cap E|$. Let $B=\{i\in[n-1-d]:|\hat{E}_{i}|<C\log{n}\}|$.
Since each $\hat{E}_{i}$ is a matching, $X_{i}$ is a binomial random variable with mean $\mathbb{E}(X_{i})=p|\hat{E}_{i}|$. For $i\not\in B$ we have $|\hat{E}_{i}|>C\log{n}$. For such $i$, by a Chernoff bound (Lemma 2.6), with probability $1-o(n^{-22})$ we have $X_{i}\geq\left(1\pm\frac{\eta}{20}\right)p^{2}|\hat{E}_{i}|$, using $p>\varepsilon\gg 1/C$ and $\eta\gg 1/C$. Thus by a union bound, with probability $1-o(n^{-21})$ we have
$|\hat{E}\cap E^{\prime}|=\sum_{i=1}^{n-d+1}X_{i}=\left(1\pm\frac{\eta}{20}\right)p^{2}|\hat{E}|+\sum_{i\in B}\left(X_{i}-\left(1\pm\frac{\eta}{20}\right)p^{2}|\hat{E}_{i}|\right).$ (57)
Let us bound the size of $B$. We have that $|\hat{E}|=\sum_{i=1}^{n-d+1}|\hat{E}_{i}|\leq|B|C\log{n}+(n-d+1-|B|)n$ and so rearranging and using (56) we obtain
$|B|\leq\frac{2n+\eta/10(n-1-d)n}{n-C\log{n}}\leq 4+\eta(n-1-d)<n^{0.99}.$
Thus, using $p>\varepsilon$ and  (56) again, we have
[…]
Now, since $0\leq X_{i}\leq|\hat{E}_{i}|$ we see that
$\left(1+\frac{\eta}{20}\right)\sum_{i\in B}|\hat{E}_{i}|\leq\sum_{i\in B}\left(X_{i}-\left(1\pm\frac{\eta}{20}\right)p^{2}|\hat{E}_{i}|\right)\leq\sum_{i\in B}|\hat{E}_{i}|,$
[…]